OpenStudy (anonymous):

Express 0.40 x 0.50 as a fraction: 1/6 Not Here 1/3 1/5 1/4

OpenStudy (anonymous):

0.4=4/10=2/5 0.5=5/10=1/2 2/5 * 1/2=2/10=1/5

OpenStudy (anonymous):

THANK YOU